Run Planning Meeting Notes, Mar 19-Mar 28, 2018

From GlueXWiki
Jump to: navigation, search

<- Back to Run Coordination Meetings:Fall2017 Spring2018 Run

Sunday, March 25 @ 10:00am

  1. Accelerator Status
    • No significant progress on increasing total beam power since yesterday
    • Started setting up Hall-B and Hall-D beams on Owl last night
    • Will work on establishing Hall-D and Hall-B beams today
      • RF expert is on site, but may not have anything new to try. May cause some loss in beam stability
    • Likely to go back to working on power issue on Monday when full compliment of experts available.
      • Beam likely to be very unstable for us then. Will know more Monday morning.
    • Beam Authorization
  2. Run Plan
    • Restablish Photon beam
    • PS High Rate Check
    • CDC Background Studies (Lubomir et al.)
    • Converter thickness study (Alex S.)
    • SSP Study (Alex S.)
    • Production running
    • Time Chart
  3. Reports
  4. Safety Issues

Monday, March 19 @ 8:45am

  1. Notes from the previous 48 hours:
    • 8:00am MCC meeting report:
      • transformers
        • 2nd Transformer installed and powered up Saturday
      • Work to restore cryo operation underway
      • Halls asked to man counting house 24/7 starting Fri. 3/23
      • Beam restoration over weekend with Physics expected to start Wed. 3/28
    • Solenoid ramped down
    • Hall D in restricted access, tagger hall
  2. Reports:
  3. AOB